OpenStudy (anonymous):

find the missing denominator for 7^16/7^12= 7^-18/?

zepdrix (zepdrix):

Hey Bey :) \[\Large\rm \frac{7^{16}}{7^{12}}=7^{\frac{-18}{x}}\]

zepdrix (zepdrix):

Applying our exponent rule on the left side of the equation, subtracting the exponents,\[\Large\rm 7^{16-12}=7^{\frac{-18}{x}}\]

zepdrix (zepdrix):

Now that the bases are the same, we can equate the exponents!\[\Large\rm 16-12=\frac{-18}{x}\]

zepdrix (zepdrix):

And solve for x.

zepdrix (zepdrix):

Or was the `?` supposed to be a denominator under the 7? Maybe I misread that :p

zepdrix (zepdrix):

Like this?\[\Large\rm \frac{7^{16}}{7^{12}}={\frac{7^{-18}}{x}}\]

OpenStudy (anonymous):

the options are 7^-6, 7^-14, 7^-22,and 7^22 so I tried solving it and didn't see my answer

OpenStudy (anonymous):

and yeah